When 50 ml of 1.5 M HCl is added to 100 ml of 1.5 M sodium hydroxide solution, temperature initially was 21.3 C and fainaly raised to 23.4C, What is kJ of heat produced? * (1 Point) 12381.86 mol 2211.05 mol 1102.9 4643.2 mo


1
Expert's answer
2021-12-15T10:03:40-0500

We use the formula

q = mc"\Delta"T


We assume that there is no heat loss to the surrounding and the resultant solution behaves as water ( Density = 1.02g/mL and c = 4.184J/g0C )

Total volume = 50mL + 100mL = 150mL

Converting to mass = 1.02g/mL x 150mL = 153g


Temperature change = 23.40C - 21.30C = 2.1oC


q = 153g x 4.184J/goC x 2.1oC

= 1317.96J

= -1317.96J ( a minus means that heat was liberated by the system)